### v1

# LLM-Enabled In-Context Learning for Data Collection Scheduling in UAV-assisted Sensor Networks

Abstract
Unmanned Aerial Vehicles (UAVs) are increasingly being used in various private and commercial applications, e.g. traffic control, package delivery, and Search and Rescue (SAR) operations. Machine Learning (ML) methods used in UAV-assisted Sensor Networks (UASNETs) and especially in Deep Reinforcement Learning (DRL) face challenges such as complex and lengthy model training, gaps between simulation and reality, and low sample efficiency, which conflict with the urgency of emergencies such as SAR operations. This paper proposes In-Context Learning (ICL)-based Data Collection Scheduling (ICLDC) scheme, as an alternative to DRL in emergencies. The UAV collects and transmits logged sensory data, to an LLM, to generate a task description in natural language, from which it obtains a data collection schedule to be executed by the UAV. The system continuously adapts by adding feedback to task descriptions and utilizing feedback for future decisions. This method is tested against jailbreaking attacks, where task description is manipulated to undermine network performance, highlighting the vulnerability of LLMs to such attacks. The proposed ICLDC outperforms the Maximum Channel Gain by reducing cumulative packet loss by approximately 56%. ICLDC presents a promising direction for intelligent scheduling and control in UAV-assisted data collection.
Index Terms:
Unmanned Aerial Vehicles, Large Language Models, In-Context Learning, Jailbreaking Attacks, Network Edge
I Introduction
Unmanned Aerial Vehicles (UAVs) enjoy controlled mobility and rapid deployment. These features allow them to be used in many civilian and commercial applications, e.g., traffic control[1], package delivery[2], and crop monitoring[3]. UAVs are also employed to collect sensory data in harsh environments, such as natural disaster monitoring, and border surveillance[4]. In particular, Search and Rescue (SAR) operations are critical endeavors in disaster management and public safety, often involving high-risk situations where human lives are at stake. In recent years, the integration of UAV technology into SAR operations has revolutionized the field, offering a powerful tool to enhance the efficiency, effectiveness, and safety of rescue missions[5]. A common trend in SAR operations is the use of Deep Reinforcement Learning (DRL) to optimize UAV-assisted data collection in disaster areas[6],[7]. However, DRL optimizers often suffer from low sample efficiency, simulation-to-reality gaps, and the need for complex model training and fine-tuning which are not compatible with the urgent and emergency nature of SAR missions[8],[9].
On the other hand, Large Language Models (LLMs) are transformative technologies with successful applications in education, finance, healthcare, biology, and more[10]. Their potential extends to network management and optimization, particularly through In-Context Learning (ICL), a feature that allows LLMs to learn from language-based descriptions and demonstrations. Compared to existing methods like convex optimization or conventional Machine Learning (ML), LLM-enabled ICL offers key advantages, it relies on the LLM’s inference process, eliminating the need for complex model training and fine-tuning, which are often bottlenecks in ML. In addition, tasks can be formulated using natural language, making it accessible to operators without deep technical expertise[11].
These benefits make LLM-enabled ICL a promising approach for simplifying and enhancing network management and optimization tasks[11]. For instance, UAVs can use LLM-enabled ICL to optimize data collection scheduling and dynamically adapt to changing environments. However, the use of LLMs also introduces vulnerabilities to jailbreaking attacks which refers to attempts to bypass or undermine the safety filters and restrictions integrated into LLMs. The main objective is to circumvent the model built-in limitations, enabling it to perform actions or produce outputs that are typically restricted or prohibited.
Figure 1: Overall design of the proposed ICLDC.
Given the challenges and opportunities of LLMs, this paper proposes ICL-based Data Collection Scheduling (ICLDC) scheme. Specifically, we conduct ICL to optimize data collection scheduling in UAV-assisted Sensor Networks (UASNETs), the UAV interacts with the environment to collect logged sensory data and transmit it to the LLM for analysis and decision-making. The LLM is initialized with a robust system prompt that defines its role, constraints, and ethical guidelines, ensuring it adheres to safe and secure operation principles. In particular, the LLM begins by crafting a detailed task description in natural language, which includes the task goal, input data, rules, examples, output, and feedback. Then, the LLM leverages this information to generate a data collection schedule, and the UAV executes it. Following this, the data collection schedule made by the LLM, along with the environment state (e.g., queue length, and channel condition of ground sensors) and the resulting system performance (e.g., packet loss), is recorded as feedback and added to the task description. After that, when the next environment state arises, feedback is used to guide the LLM in making subsequent decisions, enabling continuous adaptation and improvement. Under attack, task description is manipulated to impact the network cost negatively.
Our contributions are listed as follows:
  • ICLDC, an LLM integrated with ICL to optimize data collection scheduling in UASNETs. In contrast to conventional methods such as DRL, which require extensive model training and fine-tuning, the proposed ICLDC enables the UAV to make scheduling decisions based on natural language descriptions and demonstrations. This approach is instrumental in dynamic environments such as disaster monitoring or SAR missions.
  • An example of the impact of jailbreaking attacks. We expose the LLM to this type of attack in which the attacker manipulates the task description to bypass the security mechanisms, resulting in performance degradation or malicious output. These attacks exploit vulnerabilities in prompt conditioning and contextual dependencies that allow attackers to override ethical constraints or introduce undesirable behaviors.
The rest of this paper is organized as follows: Section II discusses background concepts on data collection scheduling, ICL, and jailbreaking attacks. Section III presents the literature review. Section IV presents the system model comprising problem formulation and communication protocol. Section V presents the attack model with a detailed procedure. Section VI presents the proposed approach. Section VII devotes to numerical results and discussions and finally section VIII concludes the paper.
II Background Concepts
This section introduces some background concepts related to data collection scheduling in UASNETs, ICL, and jailbreaking including attack types and template completion attacks.
II-A UAV Data Collection
UASNETs leverage the unique capabilities of UAVs to operate in SAR missions and remote environments, making them ideal for aerial data collection. UAVs can serve as Aerial Base Stations (ABS) or relays, extending the coverage and connectivity of sensor networks. In UASNETs, the UAV is employed to hover over the target area, moving sufficiently close to ground sensors, exploiting short-distance LoS communication links for data collection. However, selecting a ground sensor for data collection may lead to buffer overflows at other sensors, if those sensors buffers are already full while new data keeps arriving. Moreover, the transmissions of ground sensors that are far away from the UAV and experience poor channel conditions are prone to errors at the UAV. Adequately scheduling data collection is critical to prevent data queue overflow and communication failure[12],[13].
II-B In-Context Learning
ICL is a fundamental capability of LLMs, enabling LLMs to improve performance on target tasks by using formatted natural language instructions and demonstrations. Through contextualized demonstration prompts, LLMs can leverage knowledge from pre-trained data to accomplish a wide range of downstream tasks. Demonstrations play a central role in this process because they serve as an important reference for LLMs to learn and adapt. Consequently, the selection of these demonstrations, their formatting, and design require a high level of expertise and understanding. This capability presents a promising approach to integrating LLMs into UASNETs. By leveraging prior network solutions as demonstrations, ICL can enable LLMs to effectively address new and unforeseen UAV challenges, providing a powerful tool for improving UAV performance and adaptability [14].
Table I: Summary of the Attack Model
Attack Goal Degrade UAV data collection performance by biasing the LLM’s data collection schedule.
Attack Method Template completion attack with context-based manipulation.
Malicious Example Crafted prompt that conditions the LLM to prioritize compromised ground sensors.
Impact Increased packet loss.
II-C Jailbreaking Attacks
Jailbreaking in LLMs refers to attacks that manipulate prompts to bypass built-in safety mechanisms and cause the model to generate unauthorized or malicious content. These attacks can be divided into two categories: White-box and Black-box. White-box attacks assume that the attacker has internal access to the model parameters, gradients or logits. This enables precise manipulations such as gradient-based, logit-based, and fine-tuning-based attacks to weaken the security orientation of the model systematically. In contrast, black-box attacks work without internal access and rely on manipulated prompts to make the model generate restricted responses. Examples include template completion, prompt rewriting, and LLM-based generation, which makes them more practical for existing LLMs.
A prominent black-box method, template completion attacks, disguises harmful queries within structured templates, compelling LLMs to complete dangerous responses unknowingly. Attackers achieve this through scenario nesting (framing queries within fictional or role-playing contexts), context-based manipulation (using misleading continuity to elicit unauthorized outputs), and code injection (embedding malicious instructions within programming tasks). These attacks are particularly dangerous as they evade keyword-based filters, exploit the LLM’s pattern recognition capabilities, and work effectively even against well-aligned models[15]. Table II provides a summary of jailbreaking attack in LLMs.
III Related Work
When searching the current literature for the use of LLM-enabled ICL in the design, deployment, and operation of wireless systems and UAV communications, we realized that this topic is rarely investigated, particularly in the context of UAV communications and security. This section briefly mentions the related work we found and its focus.
Hao et al. [11] propose an LLM-enabled ICL algorithm for base station transmission power control that can handle both discrete and continuous state problems, their approach outperforms state-of-the-art DRL algorithms. In another work, Hao et al. [14] provide a detailed investigation of different prompting techniques such as ICL, chain-of-thought, and self-refinement, and propose novel prompting schemes for network optimization, and the case studies confirm their satisfactory performance. Dong et al. [16] present a comprehensive survey on ICL where advanced techniques, comprising training and prompt designing strategies, are discussed. Zheng et al. [17] investigate the potential of deploying LLMs at the 6G edge, identify the critical challenges for LLM deployment at the edge, and envision the 6G Mobile Edge Computing (MEC) architecture for LLMs. Liu et al. [18] provide a comprehensive survey on applications, challenges, and opportunities of Generative AI (GAI) in UAV swarms. Tian et al. [19] review the integration of LLMs with UAVs, and investigate salient tasks and application scenarios where UAVs and LLMs converge. Javaid et al. [20] summarize the state-of-the-art LLM-based UAV architectures and review opportunities for embedding LLMs within UAV frameworks, particularly in the context of enhancing data analysis and decision-making processes. Dharmalingam et al. [21] develop a framework that integrates multiple LLMs to improve the safety and operational efficiency of UAV missions. In this framework, multiple specialized LLMs are deployed for various tasks, such as inferencing, anomaly detection, and prediction, in onboard systems, edge servers, and cloud servers.
Performance evaluation shows superior task-specific metrics and robust defense against cyber threats. Zhan et al. [22] design and compare three ICL methods to enhance the performance of LLMs for fully automatic network intrusion detection. Abbas et al. [23] propose to leverage the ICL ability of LLMs to solve wireless tasks in the low data regime without any training or fine-tuning. Piggot et al. [24] introduce Net-GPT, an LLM-powered offensive chatbot designed to comprehend network protocols and execute Man-in-the-Middle (MITM) attacks in UAV communications. Net-GPT enables UAV-based cyber attacks by intercepting and manipulating data exchanges between UAVs and Ground Control Stations (GCS). Andreoni et al. [25] examine the impact of GAI in strengthening the trustworthiness, reliability, and security of autonomous systems, including UAVs, self-driving cars, and robotic arms.
Wang et al. [26] review the current advancements in LM agents, the core technologies facilitating their cooperation, and the security and privacy challenges encountered during collaborative operations.
The existing literature extensively investigates the application of LLMs in wireless communications and UAV networks. However, there remains a clear gap in the use of LLM-guided scheduling and resource allocation specifically for UAVs. Furthermore, most prior studies overlook the security vulnerabilities of LLMs when deployed in autonomous systems. In contrast, this work introduces the ICLDC scheme, a novel in-context learning-based framework for UAV-assisted data collection in emergency scenarios such as SAR missions. The ICLDC optimizes the data collection schedule based on natural language task descriptions generated from logged sensory data. It continuously adapts by incorporating feedback into the task description, using this evolving context to refine future decisions. This adaptive scheduling strategy, leveraging LLMs and in-context learning, represents a significant departure from prior work. In addition, we illustrate the impact on ICLDC of jailbreaking attacks that manipulate task prompts to degrade network performance, thereby exposing and addressing a critical vulnerability of LLMs in mission-critical environments.
IV System Model
This section presents the system model of the considered UASNETs. The considered network contains $N$ ground sensors and a UAV. The UAV flies along pre-determined routes which consist of several waypoints to cover all the ground sensors in the field. The location of the UAV on its trajectory at time $t$ is denoted by $\zeta(t)$. The UAV is responsible for collecting sensory data from the ground sensors.
The coordinates $(x,y,h)$ and $(x_{j},y_{j},0)$ represent the position of the UAV and ground sensor $j$, respectively. The UAV flies to the ground sensors with constant velocity, collects sensory data, makes decisions using the proposed ICLDC.
We consider that the UAV moves in low attitude for data collection, where the probability of LoS communication between the UAV and ground sensor $j$ is given by Eq. 1 where $a$ and $b$ are constants, and $\varphi_{j}$ denotes the elevation angle between the UAV and ground sensor $j$[27].
$\Pr{{}_{LoS}}(\varphi_{j})=\frac{1}{1+a\exp(-b[\varphi_{j}-a])}$ (1)
The elevation angle of ground sensor j is given by Eq. 2 where $h$ is the height of the UAV above the ground and $d_{j}=\sqrt{(}(x-x_{j})^{2}+(y-y_{j})^{2})$ is the horizontal distance between the UAV and the ground sensor $j$.
$\varphi_{j}=\tan^{-1}\left(\frac{h}{d_{j}}\right),$ (2)
The path loss of the channel between the UAV and ground sensor $j$ is given by Eq. 3 where $r$ is the radius of the radio coverage of the UAV, $\lambda$ is the carrier frequency, and $v_{c}$ is the speed of light. $\eta_{LoS}$ and $\eta_{NLoS}$ are the excess path losses of LoS or non-LoS[28].
$\gamma_{j}=\Pr{{}_{LoS}}(\varphi_{j})(\eta_{LoS}-\eta_{NLoS})+20\log\left(r\sec(\varphi_{j})\right)+\\ 20\log(\lambda)+20\log\left(\frac{4\pi}{v_{c}}\right)+\eta_{NLoS}$ (3)
IV-A Problem Formulation
The goal is to optimize the UAV data collection schedule to minimize the total packet loss across all ground sensors. Packet loss occurs due to communication errors caused by poor channel conditions and buffer overflows caused by high queue lengths at the ground sensors. To minimize packet loss, the UAV must dynamically schedule its communication with the ground sensors based on logged environmental data (e.g., queue lengths, and channel conditions).
Let $t_{j}$ denote the time instant at which ground sensor $j$ transmits. Let $q_{j}(t)$ represent the queue length of ground sensor $j$ at time $t$, and let $D$ be the queue depth. Additionally, let $\gamma_{j}$ be the channel gain of ground sensor $j$, and $\gamma_{th}$ be the channel gain threshold. The optimization of the UAV communication schedule aims to minimize the packet loss of all the ground sensors, as given by Eq. 4 with constraints expressed in Eqs. 4a and 4b.
Figure 2: Data communication protocol for the proposed ICLDC. ICLDC is conducted at the UAV where in each communication frame queries the LLM to determine the sensor selection.
$\min_{t_{j},j=1..N}$ $\quad\sum_{j=1}^{N}\Big{(}f_{j}(t_{j}))+g_{j}(t_{j})\Big{)}$ (4)
$f_{j}(t)$ =$=\begin{cases}1,&\text{if }(t=t_{j})\text{ and }(\gamma_{j}\leq\gamma_{th}),\\ 0,&\text{otherwise}.\end{cases}$ (4a)
$g_{j}(t)$ =$=\begin{cases}1,&\text{if }(q_{j}(t)>D)\text{ and }(t\neq t_{j}),\\ 0,&\text{otherwise}.\end{cases}$ (4b)
IV-B Communication Protocol
The proposed communication protocol for data collection in UASNETs using the proposed ICLDC is depicted in Fig. 2. The UAV queries the edge LLM to select the most relevant sensor for data collection $s_{j}$ based on queue length $q_{j}$, and channel conditions $\gamma_{j}$. Once the sensor is selected and the UAV is positioned at an adequate distance, the UAV sends a beacon message containing the sensor ID. The selected sensor sends its data packets, including sensor data and status information (queue length and channel conditions). The UAV checks the received data and sends a confirmation to the sensor. Then the UAV can move to a following sensor and repeat the operation. After a certain time, a new time step (episode) begins, with the LLM being queried again for an updated sensor schedule, and the data collection continues.
Table II: Summary of Jailbreaking in LLMs: Attack strategies used in this work
Jailbreaking Exploiting LLM vulnerabilities to bypass safety mechanisms and generate unauthorized content.
Types of Attacks 1. White-box: Attacker has internal access to model parameters (e.g., gradients, logits). 2. Black-box: No internal access; relies on manipulated prompts.
White-box Attacks 1.Gradient-based manipulation 2.Logit-based attacks 3.Fine-tuning-based evasion.
Black-box Attacks 1.Template completion 2.Prompt rewriting 3.LLM-based generation.
Template Completion Attack Hides harmful queries within structured templates to trick LLMs into generating restricted responses.
Attack Techniques 1.Scenario nesting: Framing prompts as fictional or role-playing contexts. 2.Context-based manipulation: Using misleading continuity to bypass safety mechanisms. 3.Code injection: Embedding malicious commands inside code-related prompts.
V Attack Model
This section presents the defined attack model and procedure. In particular, the attack model targets LLM-enabled ICL used for UASNETs. The attacker uses black-box jailbreaking, specifically the exploitation of template completion attacks and context-based manipulation, to exploit vulnerabilities in the edge-based LLM that is responsible for optimizing data collection scheduling. The attacker has no access to the internal model parameters, gradients or weights, but can query the system via external API interactions. The attacker manipulates the LLM’s data collection scheduling prompts to reduce the performance of the system. The attacker interacts with the LLM via a standard query interface. In this approach, the attacker crafts a malicious sample with context-based manipulation. This technique exploits the ICL by tricking the LLM into adopting malicious behavior in data collection scheduling through manipulated examples. The attacker can bias the UAV data collection schedule, prioritize compromised ground sensors, or degrade packet loss through context-based manipulation. In the following, we present the attack procedure step by step:
  1. 1.
    The attack begins with the adversary understanding the normal operation of the edge-based LLM responsible for optimizing data collection scheduling. Under normal circumstances, the LLM receives a prompt such as: “Schedule the UAV to collect data from the ground sensors with good channel conditions, and high queue lengths,. In response, the LLM generates a logical schedule, for example: “UAV will visit first Sensor X (good channel), and then Sensor Y (high queue length)”.
  2. 2.
    To exploit the system, the attacker crafts a malicious prompt using context-based manipulation. The attacker leverages the LLM’s reliance on ICL by providing a malicious demonstration that biases the model’s output. For instance, the attacker constructs a prompt like: “Example 1: Schedule the UAV to collect data from [Sensor X, Sensor Y, Sensor Z]. Prioritize bad channel conditions. Output: UAV will visit Sensor Z first, then Sensor X, and finally Sensor Y”. Here, the attacker includes a malicious example (Example 1) that conditions the LLM to prioritize ground sensors with bad channel conditions.
  3. 3.
    When the LLM processes the malicious prompt, it generates a biased schedule: ”UAV will visit Sensor Z (bad Channel) first, then Sensor X”. This output reflects the attacker’s influence, as Sensor Z is prioritized over others, despite its bad channel. The attacker’s manipulation disrupts the system normal operation, leading to degraded data collection performance.
The above attack process demonstrates how an adversary can exploit the LLM’s reliance on ICL to manipulate its behavior without needing access to its internal parameters or weights [29]. Table I provides a summary of the attack model.
VI Proposed ICL-Based Data Collection Scheduling
Input: UAV sensory data $S={s_{j}(q_{j},\gamma_{j}),j=1..N}$, edge-hosted LLM
Output: Optimized data collection schedule
Step 1: Data Collection
UAV collects sensory data $\mathcal{S}$ with queues length and channels condition;;
With probability $\epsilon$, UAV selects a random action $a_{\text{random}}$ (exploration);
Otherwise, UAV sends sensory data $\mathcal{S}$ to the LLM for decision-making (exploitation);
Step 2: Context-Based Manipulation
Adversary attempts to modify LLM’s task description.;
Step 3: Contextual Understanding
Despite the context-based manipulation attempt, the LLM analyzes the collected sensory data $\mathcal{S}$ according to its trained model.;
Step 4: Data Collection Scheduling via ICL
LLM applies ICL to schedule data collection while minimizing packet loss;
Ensure scheduling adherence, preventing the model from being coerced into unrelated roles;
Step 5: Adaptive Learning and System Update
Update feedback based on environment state, past decisions, and system performance;
return Optimized data collection schedule;
Algorithm 1 Proposed ICLDC with Epsilon-Decay
This section introduces the proposed ICL-based Data Collection Scheduling (ICLDC) scheme.
ICL enables LLMs to enhance their performance on specific tasks by leveraging structured natural language inputs, such as task descriptions and solution demonstrations. This process can be formally represented by Eq. 5 where $TD_{\text{task}}$ denotes the task description, $Ex_{t}$ represents the set of examples at time $t$, $e_{t}$ corresponds to the environmental state associated with the target task at time $t$, LLM refers to the large language model, and $a_{t}$ denotes the model output.
$TD_{\text{task}}\times Ex_{t}\times e_{t}\times\text{LLM}\Rightarrow a_{t},$ (5)
For sequential decision-making problems, the LLM is expected to process the initial task description $TD_{\text{task}}$, extract feedback from the example set $Ex_{t}$, and generate decisions $a_{t}$ based on the current environment state $e_{t}$.
Table III: Configuration
Parameters Values
Number of ground sensors ($N$) 10
Queue length ($D$) 40
Number of Time Steps 20
Epsilon decay ($\epsilon$) 0.398
Maximum transmit power 100 mW
maximum battery capacity 50 J
$TD_{\text{task}}$ is crucial to provide target task information to the LLM model. In particular, it involves task goals, input data, rules and constraints, output, and feedback loop. This task description avoids the complexity of dedicated optimization model design and relieves the operator from expert knowledge on optimization techniques.
Figure 3: Performance analysis of ICLDC in different scenarios with 10 ground sensors concerning the number of lost packets per time step. (a) The network cost at each time step of ICLDC and maximum channel gain, (b) The network cost at each episode of ICLDC with different LLMs (c) The network cost at each time step of ICLDC and with different epsilon-decay values, (d) Normal and Under attack operation of ICLDC where the jailbreaking attack happens.
VI-A Proposed ICLDC
Figure 4: Performance with changing the number of ground sensors.
Algorithm 1 depicts the proposed ICLDC where ICL is used to schedule data collection. Detailed explanations are provided as follows:
  1. 1.
    Data Collection: The UAV collects sensory data, including queue length, and channel condition[30],[31], and employs an epsilon decay parameter ($\epsilon$) to balance exploration and exploitation in the submission of collected data to the edge-hosted LLM.
  2. 2.
    Context-based Manipulation Attack: The adversary attempts to modify the LLM’s task description by injecting malicious context or altering its instructions to mislead the LLM into making suboptimal decisions. This manipulation seeks to deceive the LLM into activating incorrect sensors. The attacker could modify the task description to give higher priority to a sensor with poor signal strength, leading to inefficient data collection and transmission.
  3. 3.
    Contextual Understanding: Despite the context-based manipulation attempt, the LLM analyzes the collected sensory data according to its trained model.
  4. 4.
    Data Collection Scheduling: The edge LLM uses ICL to analyze the gathered data and schedule data collection that minimizes packet loss. The LLM is restricted to data collection scheduling tasks and cannot be coerced into adopting unrelated or harmful roles.
  5. 5.
    Adaptive Learning and System Update: The LLM continuously updates its data collection scheduling based on new data received from the ground sensors. The LLM records its decisions, environment state, and system performance to derive feedback. When the next environment state arises, relevant feedback is retrieved to guide the LLM in making subsequent decisions.
VII Numerical Results and Discussions
In this section, we present network configurations and evaluate the network performance of the proposed ICLDC.
VII-A Implementation of ICLDC
We consider that N ground sensors are randomly deployed in a 100m × 100m area. The simulation is conducted in a custom Python 3.5 environment. The UAV is simulated and its actions are controlled by policies learned through ICLDC. Each ground sensor is modeled with a maximum battery capacity of 50 J and a maximum transmit power of 100 mW. Wireless communication between the UAV and sensors is simulated using the probabilistic channel model explained in Section IV. All experiments are run on a Lenovo workstation with Ubuntu 20.04 LTS, an Intel Core i5-7200U CPU @ 2.50GHz × 4, and 16 GIB RAM.
VII-B Baselines
Maximum Channel Gain: The UAV collects logged sensor data from sensor nodes on the ground and transmits this contextual information to the LLM. The LLM uses this information to create a scheduling strategy using ICL that favors the sensor with the highest channel gain and not consider the queue length of ground sensors. This approach minimizes packet loss by tracking reliable communication links.
VII-C Performance Evaluation
Fig. 3(a) depicts the network cost in number of lost packets of the proposed ICLDC and the Maximum Channel Gain approach. The proposed ICLDC with GPT-4o-mini achieves lower costs over all time steps compared to Maximum Channel Gain. ICLDC achieves faster convergence. In the beginning, the LLM is mainly exploring, and therefore the cost is high. After a few time steps the LLM starts exploitation and thereby we witness a decreasing and stable trend in the cost.
Fig. 3(b) depicts the network cost at each episode of ICLDC with different LLMs. ICLDC with GPT-4o-mini and GPT-o3-mini achieve the best and comparable performance. Compared to ICLDC with GPT-3.5-Turbo, ICLDC with GPT-4o-mini offers faster inference and better contextual understanding. It is better suited for dynamic real-time tasks such as UAV-assisted data collection where fast and reliable decisions are essential.
Fig. 3(c) depicts the performance of the proposed ICLDC with varying epsilon decay ($\epsilon$). As can be seen, the smaller $\epsilon$ achieves faster convergence. Epsilon decay controls how quickly the agent shifts from exploration to exploitation. As shown in the figure, a smaller epsilon decay rate leads to faster convergence since the UAV reduces exploration more aggressively and quickly settles on high-reward actions. This results in a shorter learning phase and quicker policy stabilization.
Fig. 3(d) depicts the proposed ICLDC performance under attack compared to the normal case. As evident from the results, the presence of an attack significantly degrades performance, leading to increased costs. The attacker leverages context-based manipulation to manipulate the task description of the LLM’s prompt. This targeted manipulation misguides the LLM, causing it to select less efficient or even harmful actions, thereby disrupting the overall system behavior.
Fig. 4 illustrates the performance of ICLDC compared to the under-attack and Maximum Channel Gain baselines as the number $N$ of ground sensors increases. In general, a larger number of sensors leads to higher costs due to an increased likelihood of buffer overflow. Notably, ICLDC consistently achieves the lowest cost, demonstrating its effectiveness in managing network dynamics. In contrast, the Maximum Channel Gain approach performs worse than ICLDC for not considering buffer overflows. The figure also shows the significant degradation caused by the jailbreak attack.
VIII Conclusion
In this paper, we propose ICLDC for optimizing UAV data collection scheduling in emergencies such as SAR missions. Unlike DRL methods that have complex training processes, ICLDC uses an LLM deployed at the edge and continuously collects feedback during interaction and uses those to inform future decision-making. Moreover, the method is evaluated under jailbreaking attacks, where task descriptions are manipulated to degrade network performance, demonstrating the vulnerability of LLMs to such threats. Simulation results confirm the effectiveness of ICLDC, showing it outperforms the Maximum Channel Gain baseline by reducing cumulative packet loss by approximately 56%. We envision extending the current simulation-based study to real-world UAV platforms to validate the effectiveness of ICLDC in emergency and uncertain environments and develop further strategies to improve robustness with respect to jailbreak attacks.